# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

### v1

@@ E.1 Theoretical analysis
We take the derivative over $\lambda$, using that $\partial\lambda_{1}/\partial\lambda=1$ and $\partial\lambda_{2}/\partial\lambda=-1$. Let $Z(\lambda):=\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}\exp({\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}}})$. We first evaluate the derivative of this expression:
@@
Using this result we get:
@@
$=\frac{{e^{\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}})}Z(\lambda)-e^{\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}}}Z(\lambda)\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})\bm{\mathrm{\omega}}(\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}}{Z(\lambda)^{2}}$ (103)
$=\frac{{e^{\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}})}-e^{\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}}}\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})\bm{\mathrm{\omega}}(\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}}{Z(\lambda)}$ (104)
@@
For this expression we study the sign of the derivative. Notice that we can focus only on the following term since the others are always positive
@@
$(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}})-\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})\bm{\mathrm{\omega}}(\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}$ $\geq(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}})-\max_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})$ (107)
@@
$(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{2}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{2}})-\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})\bm{\mathrm{\omega}}(\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}$ $\leq(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{2}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{2}})-\min_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})$ (112)
@@
$\frac{p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})}{p(\bm{\mathrm{y}}_{1}\mid\bm{\mathrm{c}}_{1})}$ $=\frac{p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})}{p(\bm{\mathrm{y}}_{1}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})}\frac{p(\bm{\mathrm{y}}_{1}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})}{p(\bm{\mathrm{y}}_{1}\mid\bm{\mathrm{c}}_{1})}$ (117)
@@
We proceed to substituting explicitly the expression for $p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}$ into the upper bound of Eq. 118:
@@
Taking the product of Eq. 122 and Eq. 123 we obtain:
@@
$\leq M^{\lambda_{2}-\lambda_{1}}M^{\lambda_{1}-\lambda_{2}}$ (124)
$=1$ (125)
@@
$\geq\min\left(\frac{p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})^{2}}{p(\bm{\mathrm{y}}_{1}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})^{2}},\frac{p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})^{2}}{p(\bm{\mathrm{y}}_{2}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})^{2}}\right)$ (126)
@@
This last expression is in line with the condition of Assumption 3.7, showing that, for all $\bm{\mathrm{y}}\in\mathcal{Y}$, either $\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\})_{\bm{\mathrm{y}}_{1}}$ or $\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}_{2}}$ are greater or equal to $\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}$. Combining step (1) and (2), we obtain that Assumption 3.7 holds and that:
@@
We investigate experimentally whether the models used in our evaluation satisfy Assumption 3.7. We proved in Section E.1 that the DSL inference layer reduces to Probabilistic Logic methods and only consider DPL and DPL as representatives. For CBNM we conduct a separate investigation.
We first consider the case where the prior knowledge $\mathsf{K}$ is used to define the inference layer as customary in DPL. Accordingly, we found that Assumption 3.7 is satisfied for all possible couples $(\bm{\mathrm{c}}_{1},\bm{\mathrm{c}}_{2})\in[100]^{2}$ that predicts distinct labels, see Fig. 10.
We now turn to DPL. Due to the linearity of the inference layer of DPL it suffices to consider random weights as a representative for a candidate learned inference layer. Consistently, we found that for all possible couples $(\bm{\mathrm{c}}_{1},\bm{\mathrm{c}}_{2})\in[100]^{2}$ that predicts distinct labels, Assumption 3.7 is satisfied. See Fig. 11.
Finally, we evaluate explicitly whether a trained CBNM on MNIST-Add close to optimal likelihood respects, or at least is close to satisfy, Assumption 3.7. We train a CBNM with $150$ epochs of training, reaching a mean negative log-likelihood of $0.0884$ on the test set. Notice that, due to the softmax operator, there are no choices giving zero negative log-likelihood. We found that Assumption 3.7 is satisfied for the $95\%$ of possible couples $(\bm{\mathrm{c}}_{1},\bm{\mathrm{c}}_{2})\in[100]^{2}$ giving different labels. For the remaining $5\%$ the assumption is marginally violated with a maximum increase of:
$\max_{\bm{\mathrm{c}}_{1},\bm{\mathrm{c}}_{2},\lambda}\left(\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{i}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{j}\})_{\bm{\mathrm{y}}}-\max\big{(}\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{c}}_{1}\})_{\bm{\mathrm{y}}},\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}\big{)}\right)\leq 1.38\cdot 10^{-3}$ (130)
Our results on this investigation are displayed in Fig. 12.
@@
Figure 10: The logic inference layer of DPL in MNIST-Add. (Top) The behavior of $f(\lambda):=\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{i}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{j}\})$, for four pairs $i,j$ sampled randomly from the $100$ possible worlds. (Bottom) The linear layer weights for DPL.
@@
Figure 12: Optimal CBNM in MNIST-Add. We visualize the learned weights by a CBNM achieving maximum log-likelihood in MNIST-Add. (Top) The behavior of $f(\lambda):=\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{i}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{j}\})$, for four pairs $i,j$ sampled randomly the $100$ possible worlds. The sampled pairs align with Assumption 3.7. (Bottom) The linear layer weights for the trained CBNM.

### v2

@@ E.1 Theoretical analysis
We differentiating over $\lambda$, leveraging the fact that $\partial\lambda_{1}/\partial\lambda=1$ and $\partial\lambda_{2}/\partial\lambda=-1$. Let $Z(\lambda):=\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}\exp({\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}}})$. We first evaluate the derivative of this expression:
@@
This is legitimate as $Z(\lambda)>0$ by definition. Using this result we get:
@@
$=\frac{{e^{\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}})}Z(\lambda)}{Z(\lambda)^{2}}$
$\qquad-\frac{e^{\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}}}Z(\lambda)\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})\bm{\mathrm{\omega}}(\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}}{Z(\lambda)^{2}}$ (103)
$=\frac{{e^{\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}})}}{Z(\lambda)}$
$\qquad-\frac{e^{\lambda_{1}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}+\lambda_{2}\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}}}\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})\bm{\mathrm{\omega}}(\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}}{Z(\lambda)}$ (104)
@@
Next, we analyze the sign of the derivative. Notice that we can focus only on the following term since the others are always positive:
@@
$(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}})-\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})\bm{\mathrm{\omega}}(\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}$ $+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}$
$\geq(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{1}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{1}})-\max_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})$ (107)
@@
$(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{2}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{2}})-\sum_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})\bm{\mathrm{\omega}}(\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}$ $+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}$
$\leq(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}_{2}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}_{2}})-\min_{\bm{\mathrm{y}}\in\mathcal{Y}}(\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{1}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}_{2}}^{\bm{\mathrm{y}}})$ (112)
@@
$\frac{p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})}{p(\bm{\mathrm{y}}_{1}\mid\bm{\mathrm{c}}_{1})}$ (117)
$=\frac{p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})}{p(\bm{\mathrm{y}}_{1}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})}\frac{p(\bm{\mathrm{y}}_{1}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})}{p(\bm{\mathrm{y}}_{1}\mid\bm{\mathrm{c}}_{1})}$ (118)
@@
We proceed to substituting explicitly the expression for $p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}$ into the upper bound of Eq.˜119:
@@
Taking the product of Eq.˜123 and Eq.˜124 we obtain:
@@
$\qquad\leq M^{\lambda_{2}-\lambda_{1}}M^{\lambda_{1}-\lambda_{2}}$ (125)
$\qquad=1$ (126)
@@
$\hskip-185.00027pt\geq\min\left(\frac{p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})^{2}}{p(\bm{\mathrm{y}}_{1}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})^{2}},\frac{p(\bm{\mathrm{y}}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})^{2}}{p(\bm{\mathrm{y}}_{2}\mid\lambda_{1}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+\lambda_{2}\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})^{2}}\right)$ (127)
@@
This last expression is in line with the condition of Assumption˜3.8, showing that, for all $\bm{\mathrm{y}}\in\mathcal{Y}$, either $\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\})_{\bm{\mathrm{y}}_{1}}$ or $\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}_{2}}$ are greater or equal to $\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}$. Combining step (1) and (2), we obtain that Assumption˜3.8 holds and that:
@@
Finally, we investigate experimentally whether the models used in our experiments satisfy Assumption˜3.8. We have already shown in Section˜E.1 that the DSL inference layer reduces to Probabilistic Logic methods and, therefore, we only consider DPL and DPL as representatives, and our empirical results support this claim. We evaluate CBNMs separately. In this evaluation, train models on MNIST-Add, where there are two concepts $\bm{\mathrm{c}}=(c_{1},c_{2})\in[10]^{2}$ and $19$ total labels $y\in[19]$. This yields a total of $100\times 19$ weights for the inference layer $\bm{\mathrm{\omega}}$, which are fixed (by the prior knowledge) for DPL and learned from data for DPL and CBNM.
We begin from DPL, where the prior knowledge $\mathsf{K}$ defines the inference layer $\bm{\mathrm{\omega}}^{*}$. In this case, we find that – as expected – Assumption˜3.8 is satisfied by all possible pairs $(\bm{\mathrm{c}}_{1},\bm{\mathrm{c}}_{2})\in[10]^{2}\times[10]^{2}$ that predict distinct labels (as requested by the assumption), see Fig.˜11.
We now turn to DPL. Due to the linearity of the inference layer, in order to study the learned inference layer $\bm{\mathrm{\omega}}$ it suffices to consider random weights for $\bm{\mathrm{\omega}}$, cf. Eq.˜71. Also in this case, we find that for all possible pairs $(\bm{\mathrm{c}}_{1},\bm{\mathrm{c}}_{2})\in[10]^{2}\times[10]^{2}$ that predict distinct labels, Assumption˜3.8 is satisfied. See Fig.˜12.
Finally, we evaluate whether a CBNM trained on MNIST-Add and that is close to achieving optimal likelihood satisfies, at least approximately, Assumption˜3.8. To this end, we train the CBNM for $150$ epochs, reaching a mean negative log-likelihood of $0.0884$ on the test set. We find that Assumption˜3.8 is satisfied for the $95\%$ of possible pairs $(\bm{\mathrm{c}}_{1},\bm{\mathrm{c}}_{2})\in[10]^{2}\times[10]^{2}$ giving different labels. For the remaining $5\%$ the assumption is marginally violated with a maximum discrepancy:
$\max_{\bm{\mathrm{c}}_{1},\bm{\mathrm{c}}_{2},\lambda}\left(\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{i}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{j}\})_{\bm{\mathrm{y}}}-\max\big(\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{c}}_{1}\})_{\bm{\mathrm{y}}},\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}\big)\right)$ (131)
$\leq 1.38\cdot 10^{-3}$ (132)
The results are illustrated in Fig.˜13.
@@
Figure 11: The logic inference layer of DPL in MNIST-Add. (Top) The behavior of $f(\lambda):=\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{i}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{j}\})$, for four pairs $i,j$ sampled randomly from the $100$ possible worlds. The x-axis represents $\lambda$ and the y-axis $f(\lambda)$. (Bottom) The linear layer weights for DPL.
@@
Figure 13: Optimal CBNM in MNIST-Add. We visualize the learned weights by a CBNM achieving maximum log-likelihood in MNIST-Add. (Top) The behavior of $f(\lambda):=\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{i}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{j}\})$, for four pairs $i,j$ sampled randomly the $100$ possible worlds. The sampled pairs align with Assumption˜3.8. (Bottom) The linear layer weights for the trained CBNM.